Hello. May I have some assistance. I have a test and am having trouble with some. here it is. Solve by substitution method. 8x+6y=-2 and -6x+y=51 must be typed as an ordered pair. thank you
Found 3 solutions by stanbon, shree840, mananth:
Answer by stanbon(75887)

Solve by substitution method.
8x+6y=-2
-6x+y=51
-----
Solve the 2nd equation for "y":
y = 6x+51
-----
Substitute into the 1st and solve for "x":
8x+6(6x+51) = -2
8x+36x+306 = -2
44x = -308
x = -7
-----
Solve for "y":
y = 6x+51
y = 6*-7+51
y = -42+51
y = 9
-----------------
Ans: (-7,9)
===============
cheers,
Stan H.
===========

Answer by shree840(260)

8x+6y=-2
-6x+y=51
or y=51+6x substituting in the first equation above
8x+6(51+6x)=-2
8x+306+36x=-2
44x=-308
x=-308/44=-7
y=51+6x=51-42=9 ANS

Answer by mananth(16949)

8x+6y=-2 .............1
y=51+6x ---------2
/ 1 =
y = 51+6x
Plug the value of y in (1)
8x+6(51+6x)=-2
8x+306+36x)=-2
8x+36x =-2-306
44x=-308
/44
x=-7
Plug the value of x in (1)
8x+6y= -2
8*-7+6y =-2
-56+6y= -2
6y=54
/6
y=9